### v1

# Architectural Co-Design for Zero-Shot Anomaly Detection: Decoupling Representation and Dynamically Fusing Features in CLIP

Abstract
Pre-trained Vision-Language Models (VLMs) face a significant adaptation gap when applied to Zero-Shot Anomaly Detection (ZSAD), stemming from their lack of local inductive biases for dense prediction and their reliance on inflexible feature fusion paradigms. We address these limitations through an Architectural Co-Design framework that jointly refines feature representation and cross-modal fusion. Our method integrates a parameter-efficient Convolutional Low-Rank Adaptation (Conv-LoRA) adapter to inject local inductive biases for fine-grained representation, and introduces a Dynamic Fusion Gateway (DFG) that leverages visual context to adaptively modulate text prompts, enabling a powerful bidirectional fusion. Extensive experiments on diverse industrial and medical benchmarks demonstrate superior accuracy and robustness, validating that this synergistic co-design is critical for robustly adapting foundation models to dense perception tasks.
Index Terms—  anomaly detection, multimodal feature fusion, vision-language model, transfer learning, PEFT
Fig. 1: Comparison of fusion paradigms. (a) Prior works rely on a rigid, static alignment between corresponding feature blocks. (b) Our Architectural Co-Design enables a flexible fusion policy by first enriching visual features with local priors (Conv-LoRA) and then dynamically generating tailored text descriptors for each visual level (DFG).
Fig. 2: Overview of the proposed ACD-CLIP architecture. (a) The Overall Framework: Vision and text encoders are partitioned into hierarchical Groups. Each vision group is enhanced by a trainable Conv-LoRA Adapter to instill local priors. The Dynamic Fusion Gateway (DFG) then uses each visual feature $V_{i}$ to generate a tailored text descriptor for producing a level-specific anomaly map $M_{i}$. (b) The Conv-LoRA Adapter: Our parameter-efficient adapter features a multi-branch design with multi-scale convolutions inside a LoRA bottleneck.
1 Introduction
Zero-Shot Anomaly Detection (ZSAD) leverages Vision-Language Models (VLMs) [1, 2] like CLIP [3] to bypass the need for extensive, category-specific training data required by traditional methods [4, 5]. The dominant paradigm adapts these VLMs via text prompts, which has evolved from hand-crafted ensembles (e.g., WinCLIP [6]) to learnable prompts tuned on auxiliary data. These advanced methods explore object-agnostic semantics (AnomalyCLIP [7]), dynamic prompt generation (AdaCLIP [8]), and multi-layer feature queries (CLIP-AD [9]). While significantly advancing prompt engineering, this line of work faces two fundamental limitations.
First, these methods rely on an inflexible fusion paradigm [Figure 1(a)], treating the VLM’s architecture as a fixed black box. This static, layer-wise alignment assumes a fixed semantic correspondence that often fails to capture the diverse nature of anomalies. Second, a deeper representational adaptation gap exists, as CLIP’s Vision Transformer (ViT) [10] architecture lacks the local inductive biases inherent in CNNs [11, 12], which are critical for fine-grained spatial tasks.
To address these core issues, we introduce the Architectural Co-Design CLIP (ACD-CLIP) framework [Figure 1(b)], a novel approach that synergistically refines both feature representation and cross-modal fusion. The synergy is crucial: our adapter provides the fine-grained visual details that static fusion would overlook, while the gateway supplies the adaptive mechanism required to fully exploit them. Our architectural modifications are enabled by Parameter-Efficient Fine-Tuning (PEFT) [13, 14]. While recent works have shown the value of integrating convolutional structures into PEFT adapters for general vision tasks [15, 16], our work is the first to co-design such an adapter with a dynamic fusion mechanism specifically for ZSAD. Our contributions are:
  • A parameter-efficient Conv-LoRA Adapter to inject local inductive biases for fine-grained representation.
  • A Dynamic Fusion Gateway, a vision-guided mechanism enabling a flexible, bidirectional fusion policy.
  • Validation of our synergistic co-design, which yields significant performance gains on diverse ZSAD benchmarks.
2 Methodology
As illustrated in Figure 2, our ACD-CLIP framework introduces two core innovations to adapt VLMs for ZSAD: (1) an architectural co-design for decoupled representation learning, and (2) a dynamic gateway for cross-modal fusion. We detail these components below.
2.1 Hierarchical Feature Adaptation with Local Priors
To mitigate feature entanglement, we partition the vision and text encoders into $N$ sequential Groups, encouraging each to specialize in representations at a distinct semantic level. To instill local inductive priors, we integrate a parameter-efficient Conv-LoRA Adapter into each vision group. The adapter processes the sequential output from a ViT block, $X_{\text{in}}\in\mathbb{R}^{B\times L\times C}$.
The adapter features a parallel multi-branch design ($3\times 3$ and $5\times 5$ convolutions) operating within a LoRA bottleneck. This multi-scale convolutional design allows the adapter to capture local patterns at varied receptive fields, which is crucial for identifying anomalies of different sizes. For each branch $k$, the input $X_{\text{in}}$ is projected to a low-rank tensor $X_{\text{proj}}$, reshaped into a spatial format ($\mathcal{R}_{\text{2D}}$), and passed through a convolutional bottleneck comprising a strided convolution ($\text{Conv}_{\text{down}}^{(k)}$) and a transposed convolution ($\text{Conv}_{\text{up}}^{(k)}$).
$X_{\text{proj}}=X_{\text{in}}W_{\text{down}}$ (1)
$X_{\text{up}}^{(k)}=\frac{1}{k^{2}}\operatorname{Conv}_{\text{up}}^{(k)}(\operatorname{Conv}_{\text{down}}^{(k)}(\mathcal{R}_{\text{2D}}(X_{\text{proj}})))$ (2)
The output $X_{\text{up}}^{(k)}$ is reshaped back to sequence format ($\mathcal{R}_{\text{seq}}$) and projected by $W_{\text{up}}$. The outputs from all branches are then concatenated and fused by a $1\times 1$ convolution to generate the residual update $\Delta X$, which enhances the main vision stream.
$X_{\text{branch\_out}}^{(k)}=\mathcal{R}_{\text{seq}}(X_{\text{up}}^{(k)})W_{\text{up}}$ (3)
$X_{\text{cat}}=\operatorname{Concat}(X_{\text{branch\_out}}^{(3)},X_{\text{branch\_out}}^{(5)})$ (4)
$\Delta X=\mathcal{R}_{\text{seq}}(\operatorname{Conv}_{1\times 1}(\mathcal{R}_{\text{2D}}(X_{\text{cat}})))$ (5)
2.2 Dynamic Fusion Gateway
To overcome static fusion, our Dynamic Fusion Gateway (DFG) employs a vision-guided mechanism to generate custom semantic descriptors $(T_{i}^{N},T_{i}^{A})$ for each vision feature group $V_{i}$. A global context vector is extracted from $V_{i}$ and passed through an $\text{MLP}_{\text{gate}}$ to produce logits over all $N$ text feature levels. This context vector acts as a query to dynamically weigh the importance of all text levels, effectively creating a vision-guided attention mechanism over the text features. These logits are normalized via Softmax to generate dynamic weight vectors, $\bm{\omega}_{i}^{N}$ and $\bm{\omega}_{i}^{A}$:
$\bm{\omega}_{i}^{N}=[\omega_{i,1}^{N},\dots,\omega_{i,N}^{N}]=\text{Softmax}(L_{i}^{N})$ (6)
$\bm{\omega}_{i}^{A}=[\omega_{i,1}^{A},\dots,\omega_{i,N}^{A}]=\text{Softmax}(L_{i}^{A})$ (7)
These weights dynamically fuse the set of multi-level text features $\{(T_{j}^{N},T_{j}^{A})\}_{j=1}^{N}$ to produce context-aware ”normal” ($T_{i}^{N}$) and ”abnormal” ($T_{i}^{A}$) descriptors:
$T_{i}^{N}=\sum_{j=1}^{N}\omega_{i,j}^{N}\cdot T_{j}^{N}$ (8)
$T_{i}^{A}=\sum_{j=1}^{N}\omega_{i,j}^{A}\cdot T_{j}^{A}$ (9)
The final anomaly map is the average of level-specific maps $\{M_{i}\}$, each computed from the cosine similarity between $V_{i}$ and its corresponding descriptor pair.
2.3 Training Objective and Inference
The framework is trained end-to-end with a composite loss function.
Segmentation Loss. Level-specific anomaly maps $\{M_{i}\}$ are averaged to form the final prediction map $M_{\text{seg}}$, which is optimized using a combined Focal and Dice loss $\mathcal{L}_{\text{seg}}$ against the ground-truth mask $M_{\text{GT}}$:
$\mathcal{L}_{\text{seg}}=\lambda_{\text{F}}L_{\text{F}}(M_{\text{seg}},M_{\text{GT}})+\lambda_{\text{D}}L_{\text{D}}(M_{\text{seg}},M_{\text{GT}})$ (10)
Classification Loss. For image-level classification, we compute the cosine similarity between the final global visual feature ($V_{N}^{\text{cls}}$) and the original, unfused text features from the last group ($T_{N}^{N},T_{N}^{A}$), supervised via Cross-Entropy loss $\mathcal{L}_{\text{cls}}$. We use the original text features here to provide a stable, high-level semantic anchor for the classification task, decoupling it from the fine-grained fusion used for segmentation.
Total Objective. The final training objective is a weighted sum of the two losses:
$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{seg}}+\lambda_{\text{cls}}\mathcal{L}_{\text{cls}}$ (11)
where $\lambda_{\text{F}},\lambda_{\text{D}},\text{ and }\lambda_{\text{cls}}$ are balancing hyperparameters.
Inference. A single forward pass generates the anomaly map $M_{\text{seg}}$ and a classification score. The final image-level score combines the direct classification output with the maximum value from $M_{\text{seg}}$.
Table 1: Quantitative comparison with state-of-the-art methods across diverse ZSAD benchmarks. Results are reported as (AUROC, AP) in percentage. Blue: Best result. Red: Second-best result.
Domain (Metric) Dataset WinCLIP CLIP-AD AnomalyCLIP AdaCLIP ours
$N=2$ $N=3$ $N=4$ $N=6$
Industrial (Pixel-level) MVTec-AD (85.1, 18.0) (89.8, 40.0) (91.1, 34.5) (86.8, 38.1) (91.7, 44.1) (91.4, 43.6) (90.9, 44.0) (90.0, 43.1)
BTAD (71.4, 11.2) (93.1, 46.7) (93.3, 42.0) (87.7, 36.6) (96.3, 51.2) (95.9, 51.2) (96.5, 51.5) (94.6, 51.1)
MPDD (95.2, 28.1) (95.1, 28.4) (96.2, 28.9) (96.6, 29.1) (97.0, 29.8) (96.3, 30.3) (96.1, 29.4) (96.6, 30.3)
RSDD (95.1, 2.1) (99.2, 31.9) (99.1, 19.1) (99.5, 38.2) (99.1, 40.7) (99.4, 41.3) (98.9, 40.4) (98.6, 40.1)
VisA (79.6, 5.0) (95.0, 26.3) (95.4, 21.3) (95.1, 29.2) (95.7, 27.8) (95.9, 29.6) (94.6, 27.5) (94.1, 27.3)
Medical (Pixel-level) ColonDB (64.8, 14.3) (80.3, 23.7) (81.9, 31.3) (79.3, 26.2) (85.0, 35.9) (85.1, 32.6) (83.3, 35.4) (80.7, 31.1)
ClinicDB (70.7, 19.4) (85.8, 39.0) (85.9, 42.2) (84.3, 36.0) (90.4, 53.5) (89.2, 54.0) (89.8, 56.1) (85.0, 49.1)
Kvasir (69.8, 27.5) (82.5, 46.2) (81.8, 42.5) (79.4, 43.8) (88.5, 60.7) (88.8, 60.2) (88.8, 61.3) (83.3, 54.8)
BrainMRI (86.0, 49.2) (96.4, 54.2) (95.6, 53.1) (93.9, 52.3) (96.6, 55.6) (95.3, 53.0) (97.0, 61.0) (96.9, 56.1)
Liver CT (96.2, 7.2) (95.4, 7.1) (93.9, 5.7) (94.5, 5.9) (97.3, 8.8) (97.2, 7.5) (96.0, 6.8) (95.3, 6.2)
Retina OCT (80.6, 39.8) (90.9, 48.7) (92.6, 55.3) (88.5, 47.1) (96.1, 66.2) (93.7, 50.9) (91.3, 48.2) (91.5, 48.7)
Industrial (Image-level) MVTec-AD (89.3, 92.9) (89.8, 95.3) (90.3, 95.1) (90.7, 95.2) (90.9, 95.7) (90.7, 95.8) (92.4, 96.8) (90.7, 95.7)
BTAD (83.3, 84.1) (85.8, 85.2) (89.1, 91.1) (91.6, 92.4) (93.5, 96.0) (94.9, 98.0) (93.3, 94.0) (95.4, 98.2)
MPDD (63.6, 71.2) (74.5, 77.9) (73.7, 77.1) (72.1, 76.9) (78.1, 83.7) (77.6, 82.3) (74.7, 79.0) (74.8, 78.2)
RSDD (85.3, 65.3) (88.3, 73.9) (73.5, 55.0) (89.1, 70.8) (94.0, 92.9) (94.3, 92.7) (93.4, 92.2) 92.9, 91.9
VisA (78.1, 77.5) (79.8, 84.3) (82.1, 85.4) (83.0, 84.9) (85.6, 88.5) (85.5, 88.1) (83.0, 86.0) (84.1, 86.7)
Medical (Image-level) BrainMRI (82.0, 90.7) (82.8, 85.5) (86.1, 92.3) (84.9, 94.2) (89.1, 97.2) (86.8, 96.9) (88.1, 97.3) (87.3, 97.1)
Liver CT (64.2, 55.9) (62.7, 51.6) (61.6, 53.1) (64.2, 56.7) (60.2, 54.2) (65.8, 55.3) (64.4, 57.3) (68.4, 58.9)
Retina OCT (42.5, 50.9) (67.9, 71.3) (75.7, 77.4) (82.7, 80.3) (84.4, 85.6) (81.1, 80.9) (80.3, 79.1) (82.0, 79.7)
3 Experiments
We conduct extensive experiments to validate the effectiveness of our proposed Architectural Co-Design with Dynamic Fusion (ACD-CLIP). We benchmark our method against state-of-the-art (SOTA) approaches on multiple industrial and medical datasets and perform detailed ablation studies to analyze the contribution of each component.
3.1 Experimental Setup
Datasets and Baselines. We evaluate on a wide range of public benchmarks. For the industrial domain, we use MVTec-AD [17], BTAD [18], MPDD [19], RSDD [20], and VisA [21]. For the medical domain, we use BrainMRI, Liver CT, and Retina OCT from BMAD [22], ColonDB [23], ClinicDB [24], Kvasir [25] to demonstrate cross-domain generalization. We compare our method against several recent SOTA methods, including WinCLIP [6], CLIP-AD [9], AnomalyCLIP [7], and AdaCLIP [8]. To maintain a strict zero-shot setting, our model is trained on the VisA dataset for all benchmarks except VisA itself, as there is no category overlap. For evaluation on the VisA benchmark, the model is trained exclusively on the MVTec-AD dataset.
Metrics. We follow standard evaluation protocols, reporting the Area Under the Receiver Operating Characteristic curve (AUROC) and Average Precision (AP) for both Pixel-level segmentation and Image-level classification tasks.
3.2 Main Results
Our method demonstrates superior performance across diverse ZSAD benchmarks. As shown in Table 1, on the MVTec-AD benchmark, our model achieves a pixel-level AUROC/AP of 91.7%/44.0% and an image-level AUROC of 92.4%, significantly outperforming prior methods like AnomalyCLIP (91.1%/34.5% pixel; 90.3% image). This advantage is consistent across other industrial datasets such as BTAD and VisA. Crucially, our approach demonstrates strong generalization to the medical domain. For instance, on ClinicDB, our model obtains a 90.4% pixel-level AUROC, surpassing AnomalyCLIP (85.9%) by a large margin. Qualitative results (Figure 3) corroborate these findings, showing our model generates precise anomaly maps that align closely with the ground truth, exhibiting higher fidelity and less background noise than competing methods.
Fig. 3: Qualitative comparison on diverse industrial and medical datasets, showing our method’s superior localization accuracy and noise suppression over state-of-the-art baselines.
Table 2: Ablation study of our proposed components. We report the average AUROC (%) for both pixel-level and image-level anomaly detection.
Configuration Avg. AUROC
Pixel-Level Image-Level
1. Baseline 82.3 81.2
2. + Conv-LoRA Adapter 89.1 (+6.8) 84.1 (+2.9)
3. + DFG 87.6 (+5.3) 85.9 (+4.7)
4. Ours (ACD-CLIP) 91.7 90.9
3.3 Ablation and Analysis
To isolate the effects of our components, we conduct an ablation study on the MVTEC-AD dataset with a two-group configuration ($N=2$). We validate our two core contributions, the Conv-LoRA Adapter and the Dynamic Fusion Gateway (DFG), against a baseline that uses a standard linear adapter and static fusion. As shown in Table 2, both components individually yield significant improvements. Specifically, adding the Conv-LoRA Adapter boosts the pixel-level AUROC by 6.8% (to 89.1%), confirming the benefit of injecting local inductive biases for dense prediction. Concurrently, the DFG enhances the image-level AUROC by 4.7% (to 85.9%), demonstrating the effectiveness of its vision-guided, context-aware fusion. Our full model, which integrates both components, achieves the highest performance (91.7% pixel-level and 90.9% image-level AUROC), proving a strong synergistic effect between the two modules. These results validate that both components are essential and complementary for achieving a substantial performance uplift.
4 CONCLUSION
In this work, we address the core limitations of Pre-trained Vision-Language Models for Zero-Shot Anomaly Detection by introducing a novel Architectural Co-Design framework. Our approach enhances both feature representation and cross-modal fusion: by injecting local inductive biases via a parameter-efficient Conv-LoRA adapter and replacing static alignment with a vision-guided Dynamic Fusion Gateway, our model learns fine-grained, context-aware features. This synergistic design achieves highly competitive results across diverse industrial and medical benchmarks, demonstrating substantial improvements in performance and stability. Ultimately, this work validates that a co-design of representation and fusion is a critical strategy for effectively adapting foundation models to dense perception tasks. We believe this principle can be extended to other domains such as zero-shot semantic segmentation and open-vocabulary detection.