### v2

# Architectural Co-Design for Zero-Shot Anomaly Detection: Decoupling Representation and Dynamically Fusing Features in CLIP

Abstract
Pre-trained Vision-Language Models (VLMs) face a significant adaptation gap when applied to Zero-Shot Anomaly Detection (ZSAD), stemming from their lack of local inductive biases for dense prediction and their reliance on inflexible feature fusion paradigms. We address these limitations through an Architectural Co-Design framework that jointly refines feature representation and cross-modal fusion. Our method integrates a parameter-efficient Convolutional Low-Rank Adaptation (Conv-LoRA) adapter to inject local inductive biases for fine-grained representation, and introduces a Dynamic Fusion Gateway (DFG) that leverages visual context to adaptively modulate text prompts, enabling a powerful bidirectional fusion. Extensive experiments on diverse industrial and medical benchmarks demonstrate superior accuracy and robustness, validating that this synergistic co-design is critical for robustly adapting foundation models to dense perception tasks.
Index Terms—  anomaly detection, multimodal feature fusion, vision-language model, transfer learning, PEFT
Fig. 1: Comparison of fusion paradigms. (a) Prior works rely on a rigid, static alignment between corresponding feature blocks. (b) Our Architectural Co-Design enables a flexible fusion policy by first enriching visual features with local priors (Conv-LoRA) and then dynamically generating tailored text descriptors for each visual level (DFG).
Fig. 2: Overview of the proposed ACD-CLIP architecture. (a) The Overall Framework: We structure CLIP’s vision and text encoders into a hierarchy NN sequential Groups (as illustrated, $N=3$). Each vision group is enhanced by a trainable Conv-LoRA Adapter to instill local priors. The Dynamic Fusion Gateway (DFG) then uses each visual feature $V_{i}$ to generate a tailored text descriptor for producing a level-specific anomaly map $M_{i}$. (b) The Conv-LoRA Adapter: Our parameter-efficient adapter features a multi-branch design with multi-scale convolutions inside a LoRA bottleneck.
1 Introduction
Zero-Shot Anomaly Detection (ZSAD) leverages Vision-Language Models (VLMs) [1, 2] like CLIP [3] to bypass the need for extensive, category-specific training data required by traditional methods [4, 5]. The dominant paradigm adapts these VLMs via text prompts, which has evolved from hand-crafted ensembles (e.g., WinCLIP [6]) to learnable prompts tuned on auxiliary data. These advanced methods explore object-agnostic semantics (AnomalyCLIP [7]), dynamic prompt generation (AdaCLIP [8]), and multi-layer feature queries (CLIP-AD [9]). While significantly advancing prompt engineering, this line of work faces two fundamental limitations.
First, these methods rely on an inflexible fusion paradigm [Figure 1(a)], treating the VLM’s architecture as a fixed black box. This static, layer-wise alignment assumes a fixed semantic correspondence that often fails to capture the diverse nature of anomalies. Second, a deeper representational adaptation gap exists, as CLIP’s Vision Transformer (ViT) [10] architecture lacks the local inductive biases inherent in CNNs [11, 12], which are critical for fine-grained spatial tasks.
To address these core issues, we introduce the Architectural Co-Design CLIP (ACD-CLIP) framework [Figure 1(b)], a novel approach that synergistically refines both feature representation and cross-modal fusion. The synergy is crucial: our adapter provides the fine-grained visual details that static fusion would overlook, while the gateway supplies the adaptive mechanism required to fully exploit them. Our architectural modifications are enabled by Parameter-Efficient Fine-Tuning (PEFT) [13, 14]. While recent works have shown the value of integrating convolutional structures into PEFT adapters for general vision tasks [15, 16], our work is the first to co-design such an adapter with a dynamic fusion mechanism specifically for ZSAD. Our contributions are: 1) A parameter-efficient Conv-LoRA Adapter to inject local inductive biases for fine-grained representation. 2) A Dynamic Fusion Gateway, a vision-guided mechanism enabling a flexible, bidirectional fusion policy. 3) Validation of our synergistic co-design, which yields significant performance gains on diverse ZSAD benchmarks.
2 Methodology
As illustrated in Figure 2, our ACD-CLIP framework introduces two core innovations to adapt VLMs for ZSAD: (1) an architectural co-design for decoupled representation learning, and (2) a dynamic gateway for cross-modal fusion. We detail these components below.
2.1 Hierarchical Feature Adaptation with Local Priors
To mitigate feature entanglement, we partition the vision and text encoders into $N$ sequential Groups, encouraging each to specialize in representations at a distinct semantic level. To instill local inductive priors, we integrate a parameter-efficient Conv-LoRA Adapter into each vision group. The adapter processes the sequential output from a ViT block, $X_{\text{in}}\in\mathbb{R}^{B\times L\times C}$.
The adapter features a parallel multi-branch design, where each branch employs a distinct convolutional kernel size $k\times k$ for $k\in\{3,5\}$, operating within a LoRA bottleneck. This allows it to capture local patterns at varied receptive fields. The entire operation of the adapter, denoted as $\Phi_{\text{Adapter}}$, is formulated to generate a residual update $\Delta X$ as follows:
$\Delta X$ $=\Phi_{\text{Adapter}}(X_{\text{in}})$ (1)
$=\mathcal{R}_{\text{seq}}\Big{(}\operatorname{Conv}_{1\times 1}\big{(}\mathcal{R}_{\text{2D}}\big{(}\operatorname{Concat}\big{[}X_{\text{branch\_out}}^{(k)}\big{]}\big{)}\big{)}\Big{)}$
The output of each branch, $X_{\text{branch\_out}}^{(k)}$, is generated in a two-step process. First, the input tensor $X_{\text{in}}$ is passed through the LoRA down-projection and a strided down-convolution to produce an intermediate downsampled feature map, $X_{\text{down}}^{(k)}$:
$X_{\text{down}}^{(k)}=\operatorname{Conv}_{\text{down}}^{(k)}\left(\mathcal{R}_{\text{2D}}(X_{\text{in}}W_{\text{down}})\right)$ (2)
This feature map is then upsampled via a transposed convolution, reshaped, and passed through the LoRA up-projection matrix to yield the final branch output:
$X_{\text{branch\_out}}^{(k)}=\mathcal{R}_{\text{seq}}\left(\frac{1}{k^{2}}\operatorname{Conv}_{\text{up}}^{(k)}\left(X_{\text{down}}^{(k)}\right)\right)W_{\text{up}}$ (3)
Here, $W_{\text{down}}$ and $W_{\text{up}}$ are the LoRA projection matrices. $\mathcal{R}_{\text{2D}}$ and $\mathcal{R}_{\text{seq}}$ are operators for reshaping between sequence and spatial formats. $\operatorname{Conv}_{\text{down}}^{(k)}$ is a strided convolution and $\operatorname{Conv}_{\text{up}}^{(k)}$ is a transposed convolution.
2.2 Dynamic Fusion Gateway
To overcome static fusion, our Dynamic Fusion Gateway (DFG) generates custom semantic descriptors for each visual feature group $V_{i}$. A global context vector $v_{i}^{\text{global}}=\operatorname{GAP}(V_{i})$ is extracted and passed through a gating MLP to produce logits $\bm{l}_{i}^{s}$ for each semantic state $s\in\{N,A\}$ (Normal, Abnormal). These logits determine the dynamic fusion weights $\omega_{i,j}^{s}$ for the set of all text features $\{T_{j}^{s}\}_{j=1}^{N}$:
$\omega_{i,j}^{s}=\frac{\exp(l_{i,j}^{s})}{\sum_{k=1}^{N}\exp(l_{i,k}^{s})},\quad\text{where }\bm{l}_{i}^{s}=\text{MLP}_{\text{gate}}^{s}(v_{i}^{\text{global}})$ (4)
These weights dynamically fuse the multi-level text features to produce context-aware normal ($T_{i}^{N}$) and abnormal ($T_{i}^{A}$) descriptors:
$T_{i}^{s}=\sum_{j=1}^{N}\omega_{i,j}^{s}\cdot T_{j}^{s},\quad\text{for }s\in\{N,A\}$ (5)
The resulting level-specific anomaly map, $M_{i}$, is then computed by applying a softmax function over the patch-wise cosine similarities between the visual features $V_{i}$ and their corresponding dynamic text descriptors:
$M_{i}=\frac{\exp(\text{sim}(V_{i},T_{i}^{A})/\tau)}{\exp(\text{sim}(V_{i},T_{i}^{N})/\tau)+\exp(\text{sim}(V_{i},T_{i}^{A})/\tau)}$ (6)
where $\text{sim}(\cdot,\cdot)$ denotes the cosine similarity operator and $\tau$ is a temperature parameter. The final anomaly map is the average of these level-specific maps $\{M_{i}\}$.
2.3 Training Objective and Inference
The framework is trained end-to-end with a composite loss function. which jointly addresses pixel-level segmentation and image-level classification.
Segmentation Loss. The final prediction map is the average of level-specific maps, $M_{\text{seg}}=\frac{1}{N}\sum_{i=1}^{N}M_{i}$. It is optimized against the ground-truth mask $M_{\text{GT}}$ using a combined Focal [17] and Dice loss [18]:
$\mathcal{L}_{\text{seg}}=\lambda_{\text{Focal}}\mathcal{L}_{\text{Focal}}(M_{\text{seg}},M_{\text{GT}})$
$+\lambda_{\text{Dice}}\mathcal{L}_{\text{Dice}}(M_{\text{seg}},M_{\text{GT}})$ (7)
where $\lambda_{\text{Focal}}$ and $\lambda_{\text{Dice}}$ are balancing hyperparameters.
Classification Loss. For image-level classification, we compute the cosine similarity between the final global visual feature and the original, unfused text features from the last group, supervised via Cross-Entropy loss $\mathcal{L}_{\text{cls}}$. We use the original text features here to provide a stable, high-level semantic anchor for the classification task.
Total Objective. The final training objective is a weighted sum of the two losses, balanced by the hyperparameter $\lambda_{\text{cls}}$:
$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{seg}}+\lambda_{\text{cls}}\mathcal{L}_{\text{cls}}$ (8)
Inference. A single forward pass generates the anomaly map $M_{\text{seg}}$ and a classification score. The final image-level score combines the direct classification output with the maximum value from $M_{\text{seg}}$.
Table 1: Quantitative comparison with state-of-the-art methods across diverse ZSAD benchmarks. Results are reported as (AUROC, AP) in percentage. Blue: Best result. Red: Second-best result.
Domain (Metric) Dataset WinCLIP CLIP-AD AnomalyCLIP AdaCLIP ours
$N=2$ $N=3$ $N=4$ $N=6$
Industrial (Pixel-level) MVTec-AD (85.1, 18.0) (89.8, 40.0) (91.1, 34.5) (86.8, 38.1) (91.7, 44.1) (91.4, 43.6) (90.9, 44.0) (90.0, 43.1)
BTAD (71.4, 11.2) (93.1, 46.7) (93.3, 42.0) (87.7, 36.6) (96.3, 51.2) (95.9, 51.2) (96.5, 51.5) (94.6, 51.1)
MPDD (95.2, 28.1) (95.1, 28.4) (96.2, 28.9) (96.6, 29.1) (97.0, 29.8) (96.3, 30.3) (96.1, 29.4) (96.6, 30.3)
RSDD (95.1, 2.1) (99.2, 31.9) (99.1, 19.1) (99.5, 38.2) (99.1, 40.7) (99.4, 41.3) (98.9, 40.4) (98.6, 40.1)
VisA (79.6, 5.0) (95.0, 26.3) (95.4, 21.3) (95.1, 29.2) (95.7, 27.8) (95.9, 29.6) (94.6, 27.5) (94.1, 27.3)
Medical (Pixel-level) ColonDB (64.8, 14.3) (80.3, 23.7) (81.9, 31.3) (79.3, 26.2) (85.0, 35.9) (85.1, 32.6) (83.3, 35.4) (80.7, 31.1)
ClinicDB (70.7, 19.4) (85.8, 39.0) (85.9, 42.2) (84.3, 36.0) (90.4, 53.5) (89.2, 54.0) (89.8, 56.1) (85.0, 49.1)
Kvasir (69.8, 27.5) (82.5, 46.2) (81.8, 42.5) (79.4, 43.8) (88.5, 60.7) (88.8, 60.2) (88.8, 61.3) (83.3, 54.8)
BrainMRI (86.0, 49.2) (96.4, 54.2) (95.6, 53.1) (93.9, 52.3) (96.6, 55.6) (95.3, 53.0) (97.0, 61.0) (96.9, 56.1)
Liver CT (96.2, 7.2) (95.4, 7.1) (93.9, 5.7) (94.5, 5.9) (97.3, 8.8) (97.2, 7.5) (96.0, 6.8) (95.3, 6.2)
Retina OCT (80.6, 39.8) (90.9, 48.7) (92.6, 55.3) (88.5, 47.1) (96.1, 66.2) (93.7, 50.9) (91.3, 48.2) (91.5, 48.7)
Industrial (Image-level) MVTec-AD (89.3, 92.9) (89.8, 95.3) (90.3, 95.1) (90.7, 95.2) (90.9, 95.7) (90.7, 95.8) (92.4, 96.8) (90.7, 95.7)
BTAD (83.3, 84.1) (85.8, 85.2) (89.1, 91.1) (91.6, 92.4) (93.5, 96.0) (94.9, 98.0) (93.3, 94.0) (95.4, 98.2)
MPDD (63.6, 71.2) (74.5, 77.9) (73.7, 77.1) (72.1, 76.9) (78.1, 83.7) (77.6, 82.3) (74.7, 79.0) (74.8, 78.2)
RSDD (85.3, 65.3) (88.3, 73.9) (73.5, 55.0) (89.1, 70.8) (94.0, 92.9) (94.3, 92.7) (93.4, 92.2) (92.9, 91.9)
VisA (78.1, 77.5) (79.8, 84.3) (82.1, 85.4) (83.0, 84.9) (85.6, 88.5) (85.5, 88.1) (83.0, 86.0) (84.1, 86.7)
Medical (Image-level) BrainMRI (82.0, 90.7) (82.8, 85.5) (86.1, 92.3) (84.9, 94.2) (89.1, 97.2) (86.8, 96.9) (88.1, 97.3) (87.3, 97.1)
Liver CT (64.2, 55.9) (62.7, 51.6) (61.6, 53.1) (64.2, 56.7) (60.2, 54.2) (65.8, 55.3) (64.4, 57.3) (68.4, 58.9)
Retina OCT (42.5, 50.9) (67.9, 71.3) (75.7, 77.4) (82.7, 80.3) (84.4, 85.6) (81.1, 80.9) (80.3, 79.1) (82.0, 79.7)
3 Experiments
To demonstrate the efficacy and robustness of our ACD-CLIP framework, we conducted a comprehensive evaluation. This involved benchmarking our model against state-of-the-art (SOTA) competitors across a wide range of datasets and performing in-depth ablation studies to quantify the contribution of each individual component.
3.1 Experimental Setup
Datasets and Baselines. We perform a comprehensive evaluation on 12 public benchmarks spanning the industrial domain (MVTec-AD [19], VisA [20], BTAD [21], MPDD [22], RSDD [23]) and the medical domain (BrainMRI, Liver CT, and Retina OCT from BMAD [24], ColonDB [25], ClinicDB [26], Kvasir [27]). We benchmark against recent SOTA methods, including WinCLIP [6], CLIP-AD [9], AnomalyCLIP [7], and AdaCLIP [8].
Implementation Details. Our framework is built upon the pre-trained CLIP model with a ViT-L/14 backbone at a 336px resolution. To maintain a strict zero-shot setting, we train our model on an auxiliary dataset with no category overlap (VisA for all non-VisA benchmarks, and MVTec-AD for the VisA benchmark). We report the standard Area Under the Receiver Operating Characteristic curve (AUROC) and Average Precision (AP) for both pixel-level and image-level tasks.
3.2 Main Results and Analysis
As shown in Table 1, our ACD-CLIP framework consistently establishes a new state-of-the-art across both industrial and medical domains. On the widely-used MVTec-AD benchmark, our model demonstrates a clear advantage in fine-grained localization, outperforming methods like AnomalyCLIP in pixel-level AP by a significant margin of nearly +10%. This superiority in capturing precise anomaly boundaries is a direct result of our architectural co-design.
Crucially, our approach shows strong cross-domain generalization. When trained solely on industrial data, ACD-CLIP excels on challenging medical benchmarks, achieving a pixel-level AUROC of 90.4% on ClinicDB and 96.6% on BrainMRI. These quantitative gains are supported by qualitative results (Figure 3), which show our model generates significantly cleaner and more accurate anomaly maps with less background noise compared to other baselines.
Fig. 3: Qualitative comparison on diverse industrial and medical datasets, showing our method’s superior localization accuracy and noise suppression over state-of-the-art baselines.
Table 2: Ablation study of our proposed components on the MVTec-AD dataset with a two-group configuration ($N=2$). We report the average AUROC (%) for both pixel-level and image-level anomaly detection.
Configuration Avg. AUROC
Pixel-Level Image-Level
1. Baseline 82.3 81.2
2. + Conv-LoRA Adapter 89.1 (+6.8) 84.1 (+2.9)
3. + DFG 87.6 (+5.3) 85.9 (+4.7)
4. Ours (ACD-CLIP) 91.7 90.9
3.3 Ablation Study
Our ablation study (Table 2) validates the individual and synergistic contributions of our two core components against a baseline using a linear adapter and static fusion. Integrating the Conv-LoRA Adapter alone boosts the pixel-level AUROC by 6.8%, confirming its critical role in providing the local inductive biases necessary for dense prediction. Separately, adding the Dynamic Fusion Gateway (DFG) improves the image-level AUROC by 4.7%, demonstrating the effectiveness of its vision-guided, adaptive fusion mechanism. The full ACD-CLIP model, which combines both, achieves the highest performance, confirming a powerful synergistic effect between enhancing feature representation and enabling dynamic cross-modal fusion.
4 CONCLUSION
In this work, we address the core limitations of Pre-trained Vision-Language Models for Zero-Shot Anomaly Detection by introducing a novel Architectural Co-Design framework. Our approach enhances both feature representation and cross-modal fusion: by injecting local inductive biases via a parameter-efficient Conv-LoRA adapter and replacing static alignment with a vision-guided Dynamic Fusion Gateway, our model learns fine-grained, context-aware features. This synergistic design achieves highly competitive results across diverse industrial and medical benchmarks, demonstrating substantial improvements in performance and stability. Ultimately, this work validates that a co-design of representation and fusion is a critical strategy for effectively adapting foundation models to dense perception tasks. We believe this principle can be extended to other domains such as zero-shot semantic segmentation and open-vocabulary detection.